OpenStudy (anonymous):

two numbers total -8 and their difference is 6. find the numbers.

OpenStudy (johnweldon1993):

You have a little system of equations...let 'x' = number 1 and 'y' = number 2 they TOTAL -8 so x + y = -8 and they subtract to make 6 so x - y = 6 Can you solve the system?

OpenStudy (anonymous):

im stuck.

OpenStudy (johnweldon1993):

Okay well lets work through it Our 2 equations are x + y = -8 x - y = 6 Lets solve that second equation for 'x' so lets add y to both sides to get x = y + 6 Now lets go back into our first equation...and replace what we got....wherever we see an 'x' x + y = -8 We know x = y + 6 so y + 6 + y = -8 Combine like terms 2y = -14 What does y equal??

OpenStudy (anonymous):

-7

OpenStudy (anonymous):

-1 & -7

OpenStudy (johnweldon1993):

That would be correct! ...now do you understand what I did there?

OpenStudy (anonymous):

i sure do :) thanks for your help!

OpenStudy (johnweldon1993):

Awesome!...and no problem!